OpenStudy (anonymous):

The length of a rectangle is 18 m and its width is 12 m. Find the area of ½ of the rectangle.

OpenStudy (anonymous):

9*6=54 m^2

OpenStudy (anonymous):

but the answers dont match up 30 m² 108 m² 216 m² 246 m²

OpenStudy (anonymous):

its 108 m^2. My mistake. I just rushed and tried to take a shortcut

OpenStudy (anonymous):

216/2=108 54*2=108 I made a silly error and tried to to cheat but forgot to multiply by 2

OpenStudy (anonymous):

lozl thanks